OpenStudy (anonymous):

evaluate the definite integral: F 2at the top and 1at the bottom (3/x^2 - 1) dx

hartnn (hartnn):

Hi, can you do partial fractions ? or know what that is ?

OpenStudy (anonymous):

what is partial fractions?

OpenStudy (anonymous):

It is a way of "breaking apart" fractions with polynomials in them.

OpenStudy (anonymous):

split the denom as (x-1)(x+1) the the numer as 3[(x+1) - (x-1)]/2 then separate out the num... and no partial fractions will be needed

OpenStudy (anonymous):

yea we can use partial fractions

hartnn (hartnn):

something like this : \(\Large \dfrac{3}{x^2-1}=\dfrac{3}{(x+1)(x-1)}=\dfrac{A}{x-1}+\dfrac{B}{x+1}\)

OpenStudy (anonymous):

yea we can use that

hartnn (hartnn):

do you understand that ? and know how to find A,B?

OpenStudy (anonymous):

will they both be 3?

hartnn (hartnn):

not actually, you can take the LCM right ?

hartnn (hartnn):

or cross - multiply ? do it for right side

hartnn (hartnn):

to have only one common denominator of (x-1)(x+1)

OpenStudy (anonymous):

so if you cross multiply what will you get?

hartnn (hartnn):

can you try once ? show me what u get and there is error i'll correct it

hartnn (hartnn):

like a/b +c/d = (ad+bc)/bd

hartnn (hartnn):

then similarly A/(x-1)+B/(x+1) = ... ?

OpenStudy (anonymous):

i am not sure what you mean cause i have never learnt it this way so i am confused but ill send a picture what the question you asked me to do, im not sure if it is correct

hartnn (hartnn):

how did the x+1 go in the numerator ? thats incorrect...

OpenStudy (anonymous):

i'm not sure what to do im really confused right now

hartnn (hartnn):

you got this right ? a/b +c/d = (ad+bc)/bd

OpenStudy (anonymous):

yea

hartnn (hartnn):

so, using that A/(x-1)+B/(x+1) =... ?

OpenStudy (anonymous):

(a/(x+1) + (x-1)/B) / (x-1) (x+1) ?

hartnn (hartnn):

good!

hartnn (hartnn):

so, you see now that denominators are common for both sides! so we can equate the numerators we get 3 = A(x+1) +B(x-1) clear till here ?

OpenStudy (anonymous):

ohh yeaa i get that now

hartnn (hartnn):

good, so that equation is true for every value of x put x= 1 in that equation, what do u get ?

OpenStudy (anonymous):

3=A

hartnn (hartnn):

are you sure ?

hartnn (hartnn):

3 = A (1+1) A = ...?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

noo 3/2

hartnn (hartnn):

yes

hartnn (hartnn):

now to find B, you put x =-1 in that equation, what do u get ?

OpenStudy (anonymous):

-3/2

hartnn (hartnn):

yes! so did you get how we found A and B ? and \(\Large \dfrac{3}{(x+1)(x-1)}=\dfrac{A}{x-1}+\dfrac{B}{x+1}=\dfrac{3}{2}[\dfrac{1}{x-1}-\dfrac{1}{x+1}]\) ?

OpenStudy (anonymous):

yes i get that now

hartnn (hartnn):

and can you now integrate those 2 terms separately ?

OpenStudy (anonymous):

so do i now replace the x with 2 first and then take away what i get when i replace x with 1?

hartnn (hartnn):

but you need to integrate first right ? then you'll do that

hartnn (hartnn):

remember we still haven't integrated, you know the integration for 1/(x-a) general formula ?

OpenStudy (anonymous):

lnx-a right?

hartnn (hartnn):

yes! so integral of 1/(x-1) = ... ? so integral of 1/(x+1) = ... ?

OpenStudy (anonymous):

lnx-1 - lnx+1

OpenStudy (anonymous):

so the whole thing will be 3/2 x (lnx-1 - lnx+1)

OpenStudy (anonymous):

x not times

hartnn (hartnn):

yes, 3/2 (ln |x-1| - ln |x+1|) now you can do what you asked, "so do i now replace the x with 2 first and then take away what i get when i replace x with 1? "

OpenStudy (anonymous):

ohh ok i get it now thanks alot!

hartnn (hartnn):

so whats your final answer ?

OpenStudy (anonymous):

-0.61 ?

hartnn (hartnn):

no...when you put x= 1 you get ln |0| right ?

OpenStudy (anonymous):

yea so is it equal to zero?

hartnn (hartnn):

nopes, actually ln 0 = -infinity so your final answer will be +infinity got this ?

OpenStudy (anonymous):

ohh okk i see thankss again

hartnn (hartnn):

oh and since you are new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) if you have any doubts browsing this site, you can ask me. Have fun learning with Open Study! :)

OpenStudy (anonymous):

ohh okk thank you!! :)

hartnn (hartnn):

welcome ^_^